# Deep Sub-Ensembles Meets Quantile Regression: Uncertainty-aware Imputation for Time Series

### v3

@@ 5.4 Experimental Results
We evaluate imputation performance in terms of mean absolute error (MAE) and continuous ranked probability score (CRPS) [33], and the latter one is regularly adopted to evaluate probabilistic time series forecasting and measure the compatibility of an estimated probability distribution $F$ with an observation $x$. It is a common and well-defined way to evaluate uncertainty using metrics related to loss functions, for example PICP in [35] which is similar to CRPS in quantifying uncertainty. Thus, CRPS is a reasonable evaluation metric. It can be defined as:
@@
We evaluate the performance of air quality on its original data. However, since the values of different features are not in the same range, we use the normalized values to evaluate the performance of health-care. For both datasets, we report the mean and the standard error of MAE and CRPS for five trials.
Table 1: Performance comparison of CRPS for probabilistic imputation (lower is better). Bold indicates the best result and underline indicates the second-best result.
health-care air quality
10% missing 50% missing 90% missing
Multitask GP 0.489(0.005) 0.581(0.003) 0.942(0.010) 0.301(0.003)
GP-VAE 0.574(0.003) 0.774(0.004) 0.998(0.001) 0.397(0.009)
V-RIN 0.808(0.008) 0.831(0.005) 0.922(0.003) 0.526(0.025)
CSDI 0.246(0.001) 0.338(0.002) 0.528(0.002) 0.108(0.001)
CSBI 0.253(0.001) 0.352(0.002) 0.557(0.002) 0.112(0.001)
Proposed method 0.334(0.002) 0.395(0.002) 0.445(0.002) 0.162(0.001)
Table 2: Performance comparison of MAE for deterministic imputation (lower is better). Bold indicates the best result and underline indicates the second best result.
health-care air quality
10% missing 50% missing 90% missing
V-RIN 0.271(0.001) 0.365(0.002) 0.606(0.006) 25.40(0.62)
RDIS 0.319(0.002) 0.419(0.002) 0.631(0.002) 22.11(0.35)
BRITS 0.284(0.001) 0.368(0.002) 0.517(0.002) 14.11(0.26)
CSDI 0.224(0.001) 0.311(0.002) 0.484(0.003) 9.60(0.04)
CSBI 0.234(0.001) 0.315(0.002) 0.498(0.003) 9.86(0.04)
Proposed method 0.276(0.001) 0.329(0.002) 0.382(0.003) 14.05(0.04)
Table 1 presents the CRPS values for each method. As some of the baselines discussed in the paper are not specifically designed for probabilistic prediction, we have excluded them from the CRPS comparison in Table 1. Our method outperforms other baselines on both datasets except CSDI, which exhibits that the proposed method can impute missing values that largely coincide with the probability distribution of observations. In particular, our method shows the best performance and surpasses CSDI at a high missing rate (i.e., $90\%$), which suggests that it can leverage the inherent structure and correlations across time steps to impute missing values largely aligning with the distribution of observed data.
We also report the MAE of methods in Table 2. Due to the poor MAE results of the Multitask GP and GP-VAE methods, we did not list them in Table 2, which is similar to [45]. It shows that our method outperforms other baselines except CSDI and provides accurate deterministic imputations regarding imputation accuracy. Similar to CRPS, our method exhibits higher imputation accuracy than CSDI at $90\%$ missing rate.
To further explore the influence of missing rate on the proposed method and CSDI, we conducted experiments on the health-care dataset with missing rates ranging from 10% to 90% in 10% increments, as shown in Figure 4. It can be seen that compared to CSDI, our method exhibits a slower deterioration in MAE and CRPS results when the missing rate increases and is better than CSDI when the missing rate is as high as $80\%$. This shows that our methods yields a more robust imputation capability for high missing rates, which is common and more challenging in real applications.
(a) Performance of CRPS
(b) Performance of MAE
Figure 4: Performance of CRPS and MAE on health-care dataset with missing rates ranging from 10% to 90% in 10% increments.
Figure 5: Example of time series imputation for the health-care dataset with 50% missing. The red crosses denote the observed values and the blue circles denote the ground-truth. The solid and dashed lines are median values of imputations and 5% and 95% quantiles are shown as shade.
Afterwards, we demonstrate an imputation example in Figure 5. Our method demonstrates reasonable imputations with high confidence and achieves performance comparable to CSDI by effectively leveraging temporal and feature dependencies.
5.5 Ablation studies
To assess the performance with varying ensemble member counts, we also set two additional quantile levels: $Q_{2}=[0.1,0.2,...,0.9],Q_{3}=[0.05,0.10,...,0.95]$. Similarly, we train our model with 10/50/90% missing values for comparison in detail.
Fig 6 exhibits the performance of $Q_{1},Q_{2}$ and $Q_{3}$ on health-care dataset. Through analysis, it can be concluded that $Q_{1}$ with five quantile levels is sufficient to achieve the best performance whether MAE or CRPS, verifying the robustness of our method for the number of ensemble members.
(a) Performance of CRPS
(b) Performance of MAE
Figure 6: Performance of CRPS and MAE for three sets of quantile levels on health-care dataset
Table 3: Performance of CRPS and MAE for Deep Ensembles and Quantile Sub-Ensembles with quantile levels $Q_{1}$ on health-care dataset with 10% missing rate(lower is better).
10% missing 50% missing 90% missing
MAE CRPS MAE CRPS MAE CRPS
Deep Ensembles 0.286 0.402 0.358 0.496 0.407 0.556
Quantile Sub-Ensembles 0.276 0.334 0.329 0.395 0.382 0.445
Table 4: Performance comparison of MAE and CRPS for imputation about the inclusion of $\gamma$ of proposed method on health-care dataset with 10% missing rate(lower is better).
10% missing 50% missing 90% missing
MAE CRPS MAE CRPS MAE CRPS
Inclusion 0.276 0.334 0.329 0.395 0.382 0.445
Exclusion 0.301 0.413 0.352 0.468 0.396 0.497
(a) Performance of MAE
(b) Performance of CRPS
Figure 7: Performance of MAE and CRPS for the model with respect to the inclusion of L1 and L2 losses on health-care dataset with 10% missing rate.
In Algorithm 1, we employ the optimization of weighted sums of L1 and L2 losses  [10]. To elucidate their role in model training and optimization, we conducted ablation experiments to assess the impact of including L1 and L2 losses in the model. The outcomes are illustrated in Fig 7(a) and Fig 7(b). Upon meticulous analysis, it becomes apparent that the integration of L1 and L2 losses significantly enhances the training speed of the model, resulting in a more rapid convergence of results.
Table 5: Training time and inference time of the proposed method and CSDI on health-care dataset with 10% missing rate
training time(s) inference time(s)
Proposed method 621 $\pm$ 10 42$\pm$ 3
CSDI 2053$\pm$ 22 4664$\pm$ 46
@@ 5.5 Ablation studies
To evaluate the computation efficiency of the proposed method and CSDI, we conduct experiments on health-care dataset under 10% missing. In Table 5, the term “training time" refers to the time it takes for a model to go from the initial training phase to complete fitting, while “inference time" pertains to the time it takes for a trained model to perform data testing and inference. As shown in Table 5, the proposed method is significantly more computationally efficient than CSDI, and consumes much less training and inference time. Especially, CSDI needs more inference time due to performing multiple sampling regarding missing values, which hinders deployment in real-world applications. In contrast, the proposed method achieves excellent performance while reducing computation, which is beneficial to practical deployment in real-time systems.
@@ 6 Conclusion
In this paper, we introduced Quantile Sub-Ensembles, a novel uncertainty-aware approach for time series imputation that combines quantile-regression-based task networks. Our method addresses the critical issue of overconfidence in deep learning-based imputations by providing robust uncertainty estimates while maintaining high imputation accuracy, even in scenarios with high missing data rates. The ensemble framework enhances the predictive capabilities of the underlying model, offering a more reliable solution than conventional methods.
Through extensive experiments on two real-world datasets, namely health-care and air quality, our proposed method has been demonstrated its superiority over baseline approaches, particularly at higher missing rates. In comparison to state-of-the-art diffusion models like CSDI, Quantile Sub-Ensembles achieved comparable or better performance while significantly reducing computational overhead. This makes our approach not only effective but also efficient, enabling faster training and inference times.

### v4

@@ 4.4 Experimental Results
We evaluate imputation performance in terms of mean absolute error (MAE) and continuous ranked probability score (CRPS), and the latter one is regularly adopted to evaluate probabilistic time series forecasting and measure the compatibility of an estimated probability distribution $F$ with an observation $x$. It is a common and well-defined way to evaluate uncertainty using metrics related to loss functions, for example PICP in [PICP] which is similar to CRPS in quantifying uncertainty. Thus, CRPS is a reasonable evaluation metric. It can be defined as:
$CRPS(F^{-1},x)=\int_{0}^{1}2l(F^{-1}(q),x,y)dq,$ (17)
For the five datasets, we report the MAE and CRPS averaged over five trials.
Table 2: Performance comparison of proposed method and two SOTA generative models at 70% missing rate.
solar traffic healthcare
Model MAE CRPS MAE CRPS MAE CRPS
CSDI 1.137 0.187 3.382 0.083 0.372 0.406
CSBI 1.105 0.178 3.389 0.085 0.378 0.413
Proposed method 0.668 0.101 3.013 0.071 0.349 0.404
Table 3: Performance comparison of proposed method and two SOTA generative models at 90% missing rate.
solar traffic healthcare
Model MAE CRPS MAE CRPS MAE CRPS
CSDI 1.224 0.224 3.636 0.092 0.484 0.528
CSBI 1.228 0.213 3.685 0.095 0.498 0.577
Proposed method 0.766 0.112 3.228 0.075 0.372 0.441
We compare our method with baseline models on five datasets that exhibit various missing patterns. To evaluate the performance under typical missing conditions, we artificially remove 50% of the training data and using this data as ground truth. We report the results of MAE and CRPS for five trials in Table 1. It is clear that our method consistently yields the best imputation results across both CRPS and MAE metrics in most cases. These findings highlight its strength not only in capturing temporal and feature dependencies for accurate deterministic imputations, but also in delivering high-quality probabilistic estimates that closely match the distribution of observations.
To further assess the baseline methods under higher missing rates, we compare our method with two state-of-the-art generative models (CSDI and CSBI) on three datasets at 70% and 90% missing rates, as shown in Table 2 and Table 3. The results demonstrate that the proposed method surpasses both models in all experimental setups. These results emphasize that our method delivers exceptional imputation performance even when the available time series data is highly sparse, a challenging scenario often encountered in real-world applications. The performance gains are attributed to the innovative incorporation of quantile regression and sub-ensembles in our approach, which enables more accurate and reliable imputations compared to complex deep generative models.
4.5 Ablation studies
To assess the performance with varying ensemble member counts, we also set two additional quantile levels: $Q_{2}=[0.1,0.2,...,0.9],Q_{3}=[0.05,0.10,...,0.95]$. Similarly, we train our model at 50/70/90% missing values for comparison in detail.
Fig. 4 exhibits the performance of $Q_{1},Q_{2}$ and $Q_{3}$ on three datasets. Through analysis, it can be concluded that $Q_{1}$ with five quantile levels is sufficient to achieve the best performance whether MAE or CRPS, verifying the robustness of our method for the number of ensemble members.
(a) MAE of solar
(b) MAE of traffic
(c) MAE of healthcare
(d) CRPS of solar
(e) CRPS of traffic
(f) CRPS of healthcare
Figure 4: Performance of CRPS and MAE for three sets of quantile levels
Table 4: Performance of MAE and CRPS for Deep Ensembles and Quantile Sub-Ensembles with quantile levels $Q_{1}$ on three dataset at 50/70/90% missing rate.
solar traffic healthcare
Missing Model MAE CRPS MAE CRPS MAE CRPS
50% Deep Ensembles 0.807 0.102 3.124 0.082 0.339 0.397
Quantile Sub-Ensembles 0.568 0.085 2.884 0.067 0.313 0.378
70% Deep Ensembles 0.873 0.118 3.221 0.089 0.363 0.426
Quantile Sub-Ensembles 0.668 0.101 3.013 0.071 0.349 0.404
90% Deep Ensembles 0.913 0.127 3.388 0.093 0.391 0.456
Quantile Sub-Ensembles 0.766 0.112 3.228 0.075 0.372 0.441
Table 5: Performance comparison of MAE and CRPS for imputation about the inclusion and exclusion of $\gamma$ of proposed method on three dataset at 50% missing rate.
solar traffic healthcare
MAE CRPS MAE CRPS MAE CRPS
Exclusion ($\gamma$) 0.637 0.094 2.992 0.078 0.328 0.391
Inclusion ($\gamma$) 0.568 0.085 2.884 0.067 0.313 0.378
(a) MAE of solar
(b) MAE of traffic
(c) MAE of healthcare
(d) CRPS of solar
(e) CRPS of traffic
(f) CRPS of healthcare
Figure 5: Performance of MAE and CRPS for the model about the inclusion and exclusion of L1 and L2 losses on three dataset at 50% missing rate.
In Algorithm 1, we employ the optimization of weighted sums of L1 and L2 losses. To elucidate their role in model training and optimization, we conducted ablation experiments to assess the impact of including L1 and L2 losses in the model. The outcomes are illustrated in Fig. 5. Upon meticulous analysis, it becomes apparent that the integration of L1 and L2 losses significantly enhances the training speed of the model, resulting in a more rapid convergence of results.
Table 6: The comparison of training time and inference time on healthcare dataset at 50% missing rate
training time (s) inference time (s)
@@ 4.5 Ablation studies
[…]
Quantile Sub-Ensembles 621.7 $\pm$ 10.3 2.8 $\pm$ 0.3
@@
To evaluate the computational efficiency of the proposed method and measure the extent of its lead in training and inference speed, we conduct experiments on healthcare dataset at 50% missing rate with same environment. The term “training time” refers to the time it takes for a model to go from the initial training phase to complete fitting, while “inference time” pertains to the time it takes for a trained model to perform data testing and inference. As shown in Table 6, the proposed method is significantly more computationally efficient than CSDI and Deep Ensembles, consuming much less training and inference time. Especially, CSDI needs a large amount of inference time due to performing multiple sampling and complex denoising process, which hinders deployment in real-world applications. Similarly, Deep Ensembles requires training multiple independent models from scratch, leading to a substantial increase in computational cost. In contrast, the proposed method achieves excellent performance while reducing computation, which is beneficial to practical deployment in real-time systems.
(a) solar
(b) traffic
(c) healthcare
Figure 6: A case study of time series imputation for the three datasets at 50% missing rate. The red crosses represent the observed values, while the blue circles represent the ground truth. The lines represent the median values of the imputations and 5% and 95% quantiles are shown as the shade.
4.6 Case Study
To provide an intuitional comparison, we give a case study to compare the proposed method with CSBI across three datasets, as shown in Figure 6. The proposed method (green solid line) exhibits better fitting to the ground truth, learns the temporal dependency better and provides more accurate imputations compared with CSBI (gray dashed line). The area of the proposed method and CSBI between 5% and 95% quantiles are represented by the green shade and gray shade. The green shade is smaller than the gray shade, which indicates that the proposed method provides more reliable and reasonable imputation results.
@@ 5 Conclusion
In this paper, we introduced Quantile Sub-Ensembles, a novel uncertainty-aware approach for time series imputation that combines quantile-regression-based task networks. Our method addresses the critical issue of overconfidence in deep learning-based imputations by providing robust uncertainty estimates while maintaining high imputation accuracy, even in scenarios with a high missing rate. The ensemble framework enhances the predictive capabilities of the underlying model, offering a more reliable solution than conventional methods.
Through extensive experiments on five real-world datasets, our proposed method has been demonstrated its superiority over baseline approaches, particularly at higher missing rates. In comparison to state-of-the-art diffusion models like CSDI, Quantile Sub-Ensembles achieved comparable or better performance while significantly reducing computational overhead. This makes our approach not only effective but also efficient, enabling faster training and inference times.
@@
Acknowledgements
This paper is supported by the National Science Foundation of China Project (No. 62306098), Fundamental Research Funds for the Central Universities (No. JZ2024HGTB0256) and the Open Project of Anhui Provincial Key Laboratory of Multimodal Cognitive Computation, Anhui University (No. MMC202412).
Clinical trial number
Not applicable